roots of quation log2(n2-4n+5)=n-2 are


Grade:11

1 Answers

abhishek upadhyay
12 years ago

log2((n-2)2+1) = n-2

 clearly n-2 = 0 is a solution

now, (n-2)2+1 = 2n-2 has only one solution .....(by graph)

so n=2 is only solution